SWITCHlan Trouble Tickets

Transparency is very important to us. We therefore publish all trouble tickets about issues that have an impact on our customers. We show both currently open tickets as well as all closed tickets.

Ticket 20111107_1

Ticket Number: 20111107_1Ticket State: CLOSED
Ticket Opened: 2011-11-07 18:05Ticket Closed: 2011-11-07 19:00
Ticket Description: CERN-CNAF-LHCOPN-002 flapping

Problem Description:

GARR reports continuous flapping on their CERN-CNAF-LHCOPN-002 link. The link is currently unusable.
Note: there was a scheduled intervention on that DWDM node (see SWITCH ticket 20111028_1), but it was not meant to have any influence on this particular CBF link.


From 2011-11-07 11:45 until 2011-11-07 18:40
Impact: Total loss of connectivity
Service: P2P/OPN


2011-11-07 18:45
CERN and GARR reported the link as up and running.

2011-11-07 18:40
There was a patch cable which was pinched and caused an attenuation which was high enough to make the link flapping.
The cable is fixed.

2011-11-07 16:00
Our engineer is on the way back to CERN.

2011-11-07 15:00
We reached the engineer who was at CERN. He is looking into the problem as well.

2011-11-07 13:10
GARR informs about the problem. SWITCH starts to investigate. Our DWDM system does not report any problems. Testing from the Manno side reveals no indication about the nature of the problem. The engineer who was at CERN had unfortunately already left.

For all questions about this ticket, please send mail to noc@switch.ch
or call +41 44 268 15 30.